Part B The following reaction was carried out in a 3.50 L reaction vessel at 1100...

Part B

The following reaction was carried out in a 3.50 L reaction vessel at 1100 K:

C(s)+H2O(g)⇌CO(g)+H2(g)

If during the course of the reaction, the vessel is found to contain 5.50 mol of C, 15.5 mol of H2O, 4.00 mol of CO, and 8.30 mol of H2, what is the reaction quotient Q?
Answer #1

V = 3.5

K = CO*H2/(H2O)

ignore C since it is solid

[H2O] = 15.5/3.5 = 4.428

[CO] = 4/3.5 = 1.1428

[H2] = 8.3/3.5 = 2.37142

then

Q = CO*H2/H2O

Q = 1.1428*2.37142/4.428 = 0.612027727